OpenStudy (anonymous):

I need someone to show me how to solve 3x/5 -2= X+1/3. This problem is killing me.....

OpenStudy (anonymous):

(3x-10)/5=(3x+1)/3 9x-30=15x+5 4x=-35 x=-8.75